OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. 0 = x2 + 5x + 6 0 = x2 + 4x + 4 Part 2: Using complete sentences, answer the following questions about the two quadratic equations above. Do the two quadratic equations have anything in common? If so, what? What makes y = x2 + 4x + 4 different from y = x2 + 5x + 6?

OpenStudy (campbell_st):

(x +3)(x+2) = 0 then x = -2, -3 trinomial with unequal roots (x + 2)^2 = 0 then x = -2 perfect square with a repeated root The parabolas are both concave up and have a root or zero at x = -2 the perfect square on touches the x-axis once at = -2 while the trinomial cuts the x-axis at x = -3 and x = -2

OpenStudy (anonymous):

is that the answer for part1 and part 2?

OpenStudy (campbell_st):

the top is the solutions to solving the equations. the bottom... what do they have in common